# Thomason cohomology and Quillen’s Theorem A

### v1

@@ 6. Cohomology of bisimplicial sets with nontrivial coefficients
Let $X$ be a bisimplicial set and $J:\Delta(\mathrm{diag}X)\to\Delta(X)$ be the functor defined by the inclusion of simplices in the diagonal $\mathrm{diag}X$. For every coefficient system $\mathcal{M}$ for $X$, the induced homomorphism
@@
Let $X$ be a bisimplicial set. A coefficient system $\mathcal{M}$ on $\mathrm{diag}X$ is called extendable if there is a coefficient system $\mathcal{M}^{\prime}$ on $\Delta(X)$ such that $\mathcal{M}=J^{*}(\mathcal{M}^{\prime})$.
@@
$\sigma=(d_{0}\to\cdots\to d_{p})\in N\mathcal{D}_{p}\text{ and }\tau=(x_{0}\smash{\,\mathop{\longrightarrow}\limits^{\beta_{1}}\,}\cdots\smash{\,\mathop{\longrightarrow}\limits^{\beta_{q}}\,}x_{q})\in NF_{1}(d_{0})_{q},$
@@
$=\bigl{(}F_{1}(\alpha_{f_{h}(0)}\cdots\alpha_{1})(f_{v}^{*}(\tau))\bigr{)}^{\prime}.$
The second equality holds because $U(d_{0}):F_{1}(d_{0})\to F_{2}(d_{0})$ is a functor, and the third equality holds because $U$ is a natural transformation. So the equality
@@
For every $d\in\mathcal{D}$, there is a simplicial map $i_{d}:NF(d)\to\operatorname*{hocolim}_{\mathcal{D}}F$ that sends a simplex $\tau\in NF(d)_{p}$ to the pair $((d\smash{\,\mathop{\longrightarrow}\limits^{\mathrm{id}}\,}d\smash{\,\mathop{\longrightarrow}\limits^{\mathrm{id}}\,}\cdots\smash{\,\mathop{\longrightarrow}\limits^{\mathrm{id}}\,}d),\tau)$ in $N(\mathcal{D};F)_{p,p}$.
@@
Let $U:F_{1}\to F_{2}$ be a natural transformation between functors $F_{1},F_{2}:\mathcal{D}\to Cat$, and let $\mathcal{M}$ be an extendable coefficient system defined on $\operatorname*{hocolim}_{\mathcal{D}}F_{2}$. If for every $d\in\mathcal{D}$, the induced map
@@
$\mathfrak{u}^{*}:H^{*}(\operatorname*{hocolim}_{\mathcal{D}}F_{2};\mathcal{M})\to H^{*}(\operatorname*{hocolim}_{\mathcal{D}}F_{1};\mathfrak{u}^{*}\mathcal{M})$
@@
Let $\mathcal{M}^{\prime}$ be a coefficient system for the bisimplicial set $N(\mathcal{D};F_{2})$ such that $J^{*}(\mathcal{M}^{\prime})=\mathcal{M}$. By Lemma 6.5, the natural transformation $U:F_{1}\to F_{2}$ induces a bisimplicial map $\mathfrak{U}:N(\mathcal{D};F_{1})\to N(\mathcal{D};F_{2})$. Consider the following commuting diagram:
[…]
@@
$(\mathfrak{U}^{*}f)(\sigma,\tau)=(\sigma,NU(d_{0})(\tau))$
for every simplex $(\sigma,\tau)$ in $X_{1}=N(\mathcal{D};F_{1})$.
@@
is an isomorphism. The simplicial set $X_{2}^{p,*}$ is isomorphic to the disjoint sum $\coprod_{\sigma\in N\mathcal{D}_{p}}NF(d_{0})$. Hence $\mathfrak{U}^{*}$ is a disjoint union of homomorphisms
$NU(d)^{*}:H^{*}(NF_{2}(d);i_{d}^{*}J^{*}\mathcal{M}^{\prime})\to H^{*}(NF_{1}(d);i_{d}^{*}J^{*}\mathfrak{U}^{*}\mathcal{M}^{\prime}).$
By definition, we have $i_{d}^{*}J^{*}\mathcal{M}^{\prime}=i_{d}^{*}\mathcal{M}$ and $i_{d}^{*}J^{*}\mathfrak{U}^{*}\mathcal{M}^{\prime}=i_{d}^{*}\mathfrak{u}^{*}\mathcal{M}=NU(d)^{*}i_{d}^{*}\mathcal{M}$. Hence the homomorphism $NU(d)^{*}$ is an isomorphism by the assumption of the proposition. We obtain that $\mathfrak{U}^{*}$ is an isomorphism, hence $\mathfrak{u}^{*}$ is also an isomorphism. ∎
@@
[…]
@@
[…]
@@
[…]
By Theorem 1.2, the induced maps $(\kappa)^{*}$ and $(\kappa^{\prime})^{*}$ are both isomorphisms. Let
$i_{d}:N(\varphi/d)\to\operatorname*{hocolim}_{\mathcal{D}}N(\varphi/-)$
be the simplicial map which sends a simplex $\tau\in N(\varphi/d)_{p}$ to the pair $((d\smash{\,\mathop{\longrightarrow}\limits^{\mathrm{id}}\,}\cdots\smash{\,\mathop{\longrightarrow}\limits^{\mathrm{id}}\,}d),\tau)$ in $N(\mathcal{D};\varphi/-)_{p,p}$. Similarly let $i_{d}^{\prime}:N(\mathrm{id}/d)\to\operatorname*{hocolim}_{\mathcal{D}}N(\mathrm{id}/-)$ be the simplicial map defined in a similar way for the identity functor. By Proposition 6.6, $\mathfrak{u}^{*}$ is an isomorphism if for every $d\in\mathcal{D}$, the induced map
@@
is an isomorphism. We have $(i_{d}^{\prime})^{*}(\kappa^{\prime})^{*}\mathcal{M}=(j_{d}^{\prime})^{*}\mathcal{M}=\mathcal{M}_{d}$ and
$NU(d)^{*}(i_{d}^{\prime})^{*}(\kappa^{\prime})^{*}\mathcal{M}=NU(d)^{*}\mathcal{M}_{d}=\mathcal{M}_{d}^{\prime}.$
Hence $NU(d)^{*}$ is an isomorphism by the assumption of the theorem. We conclude that $u^{*}$ is an isomorphism, hence $(N\varphi)^{*}$ is an isomorphism. ∎
@@ 7. Thomason cohomology of the Grothendieck construction
Let $F:\mathcal{D}\to Cat$ be a functor. The Grothendieck construction of $F$ is the category $\int_{\mathcal{D}}F$ whose objects are the pairs $(d,x)$ with $x\in F(d)$, and whose morphisms $(d,x)\to(d^{\prime},x^{\prime})$ are given by pairs $(\alpha,\gamma)$ where $\alpha:d\to d^{\prime}$ is a morphism in $\mathcal{D}$, and $\gamma:F(\alpha)(d)\to d^{\prime}$ is a morphism in $F(d^{\prime})$.
@@
[…]
@@
where $N\mathcal{C}$ is considered as a bisimplicial map that is constant in the horizontal direction. The $(p,q)$-simplices of $N(\mathcal{D},N\widetilde{F})$ are the triples $(\sigma,\tau,\mu)$ where
@@
${\Delta(\operatorname*{hocolim}_{\mathcal{D}}N\widetilde{F}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces)}$$\scriptstyle{J}$$\scriptstyle{\lambda_{2}}$${\Delta(N\mathcal{C})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{J}$${\Delta(N(\mathcal{D};N\widetilde{F})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\lambda_{2}^{\prime}}$${\Delta(N_{b}\mathcal{C})}$
@@
[…]
@@
Consider the bisimplicial set $X=N(\mathcal{D};N\widetilde{F})$. Let $\mathcal{M}$ be a coefficient system on $N\mathcal{C}$, and let $\mathcal{M}^{\prime}=(\lambda_{2}^{\prime})^{*}\mathcal{M}$ be the coefficient system on $X$ induced by the bisimplicial map $\lambda_{2}^{\prime}:\Delta(X)\to\Delta(N_{b}\mathcal{C})$ introduced above.
For a bisimplicial set $X$ and a coefficient system $\mathcal{M}$, the cohomology $H^{*}(X;\mathcal{M})$ is defined as the cohomology of the cochain complex $C^{*}(X;\mathcal{M})$ where for each $n\geq 0$,
@@
For $X=N(\mathcal{D};N\widetilde{F})$, the vertical simplicial set $X^{p,*}$ is isomorphic to the disjoint union $\coprod_{\sigma\in N\mathcal{D}_{p}}\{\sigma\}\times N\widetilde{F}(d_{0})$. For each $\sigma\in N\mathcal{D}_{p}$, $\{\sigma\}\times N\widetilde{F}(d_{0})$ is isomorphic to the simplicial set $\{(d_{0}\to\cdots\to d_{0})\}\times N\widetilde{F}(d_{0})$. For each $d\in\mathcal{D}$, let
$i_{d}:N\widetilde{F}(d)\to\operatorname*{hocolim}_{\mathcal{D}}N\widetilde{F}$
be the simplicial map which sends a simplex $\widetilde{\tau}=(\tau,\mu)\in N\widetilde{F}(d)_{n}$ to the simplex
$((d\smash{\,\mathop{\longrightarrow}\limits^{\mathrm{id}}\,}\cdots\smash{\,\mathop{\longrightarrow}\limits^{\mathrm{id}}\,}d),\tau,\mu)$
in $N(\mathcal{D};N\widetilde{F})_{n,n}$. Since $X^{p,*}$ is a disjoint union, there is an isomorphism
$H^{q}_{v}(C^{p,*})\cong\prod_{\sigma=(d_{0}\to\cdots\to d_{p})\in N\mathcal{D}_{p}}H^{q}(N\widetilde{F}(d_{0});i_{d_{0}}^{*}J^{*}\mathcal{M}^{\prime}).$
@@
$H^{q}_{v}(C^{p,*})\cong\prod_{\sigma=(d_{0}\to\cdots\to d_{p})\in N\mathcal{D}_{p}}H^{q}(N\widetilde{F}(d_{0});j_{d_{0}}^{*}\mathcal{M}).$
The assignment $\sigma\to H^{q}(N\widetilde{F}(d_{0});j_{d_{0}}^{*}\mathcal{M})$ defines a functor $\Delta(N\mathcal{D})\to R$-Mod, hence it defines a coefficient system for $N\mathcal{D}$. We denote this coefficient system by $\mathcal{H}^{q}_{\mathcal{M}}$. The horizontal boundary maps $H_{h}^{p}(H_{v}^{q}(C^{*,*}))$ coincide with the boundary maps of the cochain complex $C^{*}(N\mathcal{D};\mathcal{H}^{q}_{\mathcal{M}})$, hence $E_{2}^{p,q}\cong H^{p}(N\mathcal{D};\mathcal{H}^{q}_{\mathcal{M}})$. The spectral sequence converges to the cohomology of the bisimplicial set $N(\mathcal{D};N\widetilde{F})$ with coefficients in $\mathcal{M}^{\prime}=(\lambda_{2}^{\prime})^{*}\mathcal{M}$. By Propositions 7.4 and 6.3, this cohomology group is isomorphic to $H^{*}_{Th}(\mathcal{C};\mathcal{M})$. ∎
@@
As a consequence of Thomason’s homotopy colimit theorem the nerve of the category $\Delta(X)_{G}$ is homotopy equivalent to
@@
The category $\Delta_{G}(X)$ is often used in the definition of $G$-equivariant cohomology groups $H^{*}_{G}(X;\mathcal{M})$ with coefficients in $\mathcal{M}:\Delta(X)_{G}\to R$-Mod. The $G$-equivariant cohomology groups of $X$ are defined as the cohomology groups of the cochain complex $C^{*}(X;\mathcal{M})$ where, for each $n\geq 0$, the $n$-th chain module is defined by

### v2

@@ 6. Cohomology of bisimplicial sets with nontrivial coefficients
Let $X$ be a bisimplicial set and $J:\Delta(\mathrm{diag}X)\to\Delta(X)$ be the functor defined by the inclusion of simplices in $\mathrm{diag}X$. For every coefficient system $\mathcal{M}$ for $X$, the induced homomorphism
@@
Let $X$ be a bisimplicial set. A coefficient system on $X$ is called horizontally constant if for every $f:[p^{\prime}]\to[p]$ and every $x\in X_{p,q}$, the induced map
$\mathcal{M}(f\times\mathrm{id}):\mathcal{M}(x)\to\mathcal{M}((f\times\mathrm{id})^{*}(x))$
is the identity map. A coefficient system $\mathcal{M}$ on $\mathrm{diag}X$ is called strongly extendable if there is a horizontally constant coefficient system $\mathcal{M}^{\prime}$ on $\Delta(X)$ such that $\mathcal{M}=J^{*}(\mathcal{M}^{\prime})$.
We have the following observation.
Lemma 6.5.
Let $\varphi:\mathcal{C}\to\mathcal{D}$ be a functor and $\kappa:X=\operatorname*{hocolim}_{\mathcal{D}}N(\varphi/-)\to N\mathcal{C}$ be the simplicial map defined in Definition 1.1. Then for every coefficient system $\mathcal{M}$ for $N\mathcal{C}$, the induced system $\kappa^{*}\mathcal{M}$ is a strongly extendable coefficient system for $X$.
Proof.
By Lemma 5.1, the simplicial map $\kappa$ is the restriction of a bisimplicial map defined by $\kappa(\sigma,\tau,\mu)=\tau$. For every $f:[p^{\prime}]\to[p]$, we have
$(\kappa^{*}\mathcal{M})(f\times\mathrm{id})=\mathcal{M}(\kappa((f\times\mathrm{id})^{*}))=\mathcal{M}(\mathrm{id})=\mathrm{id}$
by Equation 3. This shows that the bisimplicial module $\kappa^{*}M$ defined on $N(\mathcal{D},N(\varphi/-))$ is horizontally constant. Hence $\kappa^{*}\mathcal{M}$ is strongly extendable. ∎
@@
$\sigma=(d_{0}\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{1}}\,}\cdots\smash{\,\mathop{\longrightarrow}\limits^{\alpha_{p}}\,}d_{p})\in N\mathcal{D}_{p}\text{ and }\tau=(x_{0}\smash{\,\mathop{\longrightarrow}\limits^{\beta_{1}}\,}\cdots\smash{\,\mathop{\longrightarrow}\limits^{\beta_{q}}\,}x_{q})\in NF_{1}(d_{0})_{q},$
@@
$=\bigl(F_{1}(\alpha_{f_{h}(0)}\cdots\alpha_{1})(f_{v}^{*}(\tau))\bigr)^{\prime}$
where the second equality holds because $U(d_{0}):F_{1}(d_{0})\to F_{2}(d_{0})$ is a functor, and the third equality holds because $U$ is a natural transformation. So the equality
@@
For every $d\in\mathcal{D}$, there is a simplicial map $i_{d}:NF(d)\to\operatorname*{hocolim}_{\mathcal{D}}F$ that sends a simplex $\tau\in NF(d)_{p}$ to the pair $((d\smash{\,\mathop{\longrightarrow}\limits^{\mathrm{id}}\,}d\smash{\,\mathop{\longrightarrow}\limits^{\mathrm{id}}\,}\cdots\smash{\,\mathop{\longrightarrow}\limits^{\mathrm{id}}\,}d),\tau)$ in $N(\mathcal{D};NF)_{p,p}$.
@@
Let $U:F_{1}\to F_{2}$ be a natural transformation between functors $F_{1},F_{2}:\mathcal{D}\to Cat$, and let $\mathcal{M}$ be a strongly extendable coefficient system defined on $\operatorname*{hocolim}_{\mathcal{D}}F_{2}$. If for every $d\in\mathcal{D}$, the induced map
@@
$\mathfrak{u}^{*}:H^{*}(\operatorname*{hocolim}_{\mathcal{D}}NF_{2};\mathcal{M})\to H^{*}(\operatorname*{hocolim}_{\mathcal{D}}NF_{1};\mathfrak{u}^{*}\mathcal{M})$
@@
Let $\mathcal{M}^{\prime}$ be a horizontally constant coefficient system for the bisimplicial set $N(\mathcal{D};NF_{2})$ such that $J^{*}(\mathcal{M}^{\prime})=\mathcal{M}$. By Lemma 6.6, the natural transformation $U:F_{1}\to F_{2}$ induces a bisimplicial map $\mathfrak{U}:N(\mathcal{D};NF_{1})\to N(\mathcal{D};NF_{2})$. Consider the following commuting diagram:
${H^{*}(N(\mathcal{D};NF_{2});\mathcal{M}^{\prime})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathfrak{U}^{*}}$$\scriptstyle{J^{*}}$${H^{*}(N(\mathcal{D};NF_{1});\mathfrak{U}^{*}\mathcal{M}^{\prime})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{J^{*}}$${H^{*}(\operatorname*{hocolim}_{\mathcal{D}}NF_{2};\mathcal{M})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathfrak{u}^{*}}$${H^{*}(\operatorname*{hocolim}_{\mathcal{D}}NF_{1};\mathfrak{u}^{*}\mathcal{M})}$
@@
$(\mathfrak{U}^{*}f)(\sigma,\tau)=f(\sigma,NU(d_{0})(\tau))$
for every simplex $(\sigma,\tau)$ in $X_{1}=N(\mathcal{D};NF_{1})$.
@@
is an isomorphism. The simplicial set $X_{2}^{p,*}$ is isomorphic to the disjoint union $\coprod_{\sigma\in N\mathcal{D}_{p}}NF_{2}(d_{0})$ by Equation 1. Hence $H^{*}_{v}(X_{2}^{p,*};\mathcal{M}^{\prime})$ is isomorphic to the direct product
$\prod_{\sigma\in N\mathcal{D}_{p}}H^{*}(NF_{2}(d_{0}),i_{\sigma}^{*}\mathcal{M}^{\prime})$
where $i_{\sigma}:NF_{2}(d_{0})\to X_{2}^{p,*}$ is the simplicial map that sends $\tau$ to $(\sigma,\tau).$ Since $\mathcal{M}^{\prime}$ is horizontally constant, $i_{\sigma}^{*}\mathcal{M}^{\prime}=i_{d_{0}}^{*}\mathcal{M}$ as coefficients systems for $NF_{2}(d_{0})$. Hence $\mathfrak{U}^{*}$ is a disjoint union of homomorphisms
$NU(d_{0})^{*}:H^{*}(NF_{2}(d_{0});i_{d_{0}}^{*}\mathcal{M})\to H^{*}(NF_{1}(d_{0});NU(d_{0})^{*}i_{d_{0}}^{*}\mathcal{M})$
over all $\sigma:(d_{0}\to\cdots\to d_{p})$ in $N\mathcal{D}_{p}$. Each of these homomorphisms are isomorphism by the assumption of the proposition. This gives that $\mathfrak{U}^{*}$ is an isomorphism, hence $\mathfrak{u}^{*}$ is also an isomorphism. ∎
@@
${N(\varphi/d)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{j_{d}}$$\scriptstyle{NU(d)}$${N\mathcal{C}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{N\varphi}$${N(\mathrm{id}/d)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{j_{d}^{\prime}}$${N\mathcal{D}}$
@@
${\operatorname*{hocolim}_{\mathcal{D}}N(\varphi/-)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathfrak{u}}$$\scriptstyle{\kappa}$${N\mathcal{C}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{N\varphi}$${\operatorname*{hocolim}_{\mathcal{D}}N(\mathrm{id}/-)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\kappa^{\prime}}$${N\mathcal{D}}$
@@
${H^{*}(\operatorname*{hocolim}_{\mathcal{D}}N(\varphi/-);\kappa^{*}(N\varphi)^{*}\mathcal{M})}$${H^{*}(N\mathcal{C};(N\varphi)^{*}\mathcal{M})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\kappa^{*}}$${H^{*}(\operatorname*{hocolim}_{\mathcal{D}}N(\mathrm{id}/-);(\kappa^{\prime})^{*}\mathcal{M})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathfrak{u}^{*}}$${H^{*}(N\mathcal{D};\mathcal{M})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces.}$$\scriptstyle{(\kappa^{\prime})^{*}}$$\scriptstyle{(N\varphi)^{*}}$
By Theorem 1.2, the induced maps $\kappa^{*}$ and $(\kappa^{\prime})^{*}$ are both isomorphisms. We claim that $\mathfrak{u}^{*}$ is also an isomorphism. This will imply that $(N\varphi)^{*}$ is an isomorphism and complete the proof.
To see that $\mathfrak{u}^{*}$ is an isomorphism, we will apply Proposition 6.7. Note that by Lemma 6.5, for any coefficient system $\mathcal{M}$ for $N\mathcal{D}$, the induced coefficient system $(\kappa^{\prime})^{*}\mathcal{M}$ is strongly extendable. Let $i_{d}:N(\varphi/d)\to\operatorname*{hocolim}_{\mathcal{D}}N(\varphi/-)$ be the simplicial map which sends a simplex $(\tau,\mu)\in N(\varphi/d)_{q}$ to the pair $((d\smash{\,\mathop{\longrightarrow}\limits^{\mathrm{id}}\,}\cdots\smash{\,\mathop{\longrightarrow}\limits^{\mathrm{id}}\,}d),\tau,\mu)$ in $N(\mathcal{D};N(\varphi/-))_{q,q}$, and let $i_{d}^{\prime}:N(\mathrm{id}/d)\to\operatorname*{hocolim}_{\mathcal{D}}N(\mathrm{id}/-)$ be the simplicial map defined in a similar way for the identity functor. We have
$(i_{d}^{\prime})^{*}(\kappa^{\prime})^{*}\mathcal{M}=(j_{d}^{\prime})^{*}\mathcal{M}=\mathcal{M}_{d}$
and
$i_{d}^{*}\kappa^{*}(N\varphi)^{*}\mathcal{M}=i_{d}^{*}\mathfrak{u}^{*}(\kappa^{\prime})^{*}\mathcal{M}=NU(d)^{*}(i_{d}^{\prime})^{*}(\kappa^{\prime})^{*}\mathcal{M}=NU(d)^{*}\mathcal{M}_{d}.$
This gives that for every $d\in\mathcal{D}$,
@@
is an isomorphism by the assumption of the theorem. Hence we can apply Proposition 6.7 to the natural transformation $U:\varphi/-\to\mathrm{id}/$-, and conclude that $\mathfrak{u}^{*}$ is an isomorphism. We conclude that $(N\varphi)^{*}$ is an isomorphism. ∎
@@ 7. Thomason cohomology of the Grothendieck construction
Let $F:\mathcal{D}\to Cat$ be a functor. The Grothendieck construction of $F$ is the category $\int_{\mathcal{D}}F$ whose objects are the pairs $(d,x)$ with $x\in F(d)$, and whose morphisms $(d,x)\to(d^{\prime},x^{\prime})$ are given by pairs $(\alpha,\gamma)$ where $\alpha:d\to d^{\prime}$ is a morphism in $\mathcal{D}$, and $\gamma:F(\alpha)(x)\to x^{\prime}$ is a morphism in $F(d^{\prime})$.
@@
$\operatorname*{hocolim}\limits_{\mathcal{D}}NF$$\operatorname*{hocolim}\limits_{\mathcal{D}}N\widetilde{F}$$N\mathcal{C}.$$\scriptstyle{\eta}$$\scriptstyle{\lambda_{1}}$$\scriptstyle{\lambda_{2}}$
@@
[…]
@@
${\Delta(\operatorname*{hocolim}_{\mathcal{D}}N\widetilde{F}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces)}$$\scriptstyle{J}$$\scriptstyle{\lambda_{2}}$${\Delta(N\mathcal{C})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{J}$${\Delta(N(\mathcal{D};N\widetilde{F}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\lambda_{2}^{\prime}}$${\Delta(N_{b}\mathcal{C})}$
@@
${H^{*}(\operatorname*{hocolim}_{\mathcal{D}}N\widetilde{F};\lambda_{2}^{*}\mathcal{M})}$${H^{*}_{Th}(\mathcal{C};\mathcal{M})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\lambda_{2}^{*}}$${H^{*}(N(\mathcal{D};N\widetilde{F});(\lambda_{2}^{\prime})^{*}\mathcal{M})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{J^{*}}$${\ignorespaces\ignorespaces\ignorespaces\ignorespaces H^{*}(\Delta(N_{b}\mathcal{C});\mathcal{M})\ignorespaces\ignorespaces\ignorespaces\ignorespaces.}$$\scriptstyle{(\lambda_{2}^{\prime})^{*}}$$\scriptstyle{J^{*}}$
@@
[…]
For a bisimplicial set $X$ and a coefficient system $\mathcal{M}$ for $X$, the cohomology $H^{*}(X;\mathcal{M})$ is defined as the cohomology of the cochain complex $C^{*}(X;\mathcal{M})$ where for each $n\geq 0$,
@@
[…]
$\prod_{\sigma\in N\mathcal{D}_{p}}H^{q}(N\widetilde{F}(d_{0}),i_{\sigma}^{*}\mathcal{M}^{\prime}).$
Since $\mathcal{M}^{\prime}$ is horizontally constant, $i_{\sigma}^{*}\mathcal{M}^{\prime}=i_{d_{0}}^{*}J^{*}\mathcal{M}^{\prime}$ where
$i_{d_{0}}:N\widetilde{F}(d_{0})\to\operatorname*{hocolim}_{\mathcal{D}}N\widetilde{F}$
[…]
$H^{q}_{v}(C^{p,*})\cong\prod_{\sigma\in N\mathcal{D}_{p}}H^{q}(N\widetilde{F}(d_{0});i_{d_{0}}^{*}J^{*}\mathcal{M}^{\prime}).$
@@
$H^{q}_{v}(C^{p,*})\cong\prod_{\sigma\in N\mathcal{D}_{p}}H^{q}(N\widetilde{F}(d_{0});j_{d_{0}}^{*}\mathcal{M}).$
[…]
@@
As a consequence of Thomason’s homotopy colimit theorem, the nerve of the category $\Delta(X)_{G}$ is homotopy equivalent to
@@
The category $\Delta(X)_{G}$ is often used in the definition of $G$-equivariant cohomology groups $H^{*}_{G}(X;\mathcal{M})$ with coefficients in $\mathcal{M}:\Delta(X)_{G}\to R$-Mod. The $G$-equivariant cohomology groups of $X$ are defined as the cohomology groups of the cochain complex $C^{*}(X;\mathcal{M})$ where, for each $n\geq 0$, the $n$-th chain module is defined by
@@
Acknowledgements and Funding Declaration
The first author is supported by TÜBİTAK 2211-A National PhD Scholarship Program. The second author is supported by TÜBİTAK 2219-International Postdoctoral Research Fellowship Program (2023, 2nd term). We thank TÜBİTAK for their support of this research.
We also thank Matthew Gelvin and Caroline Yalçın for reading an earlier version of the paper and for their suggestions. We thank the referee for careful reading of the paper and pointing out a mistake in the earlier version of the paper.
Author contributions statement
All authors contributed to the study conception and design. The first draft of the manuscript was written by both authors and all authors commented on previous versions of the manuscript. All authors read and approved the final manuscript.
Data Availability
This manuscript does not report data generation or analysis.
Conflict of interest
The authors declare that they have no conflicts of interest.